Change the following base 2 numbers to their decimal (base 10) equivalents.
1010 ______ 1 _____ 1011101 ________ 1010111 _______

Answer by stanbon(57347)

Change the following base 2 numbers to their decimal (base 10) equivalents.
1010 = 2+8=10 1011101 = 1+4+8+16+64 = 93 1010111 = 1+2+4+16+64 = 87_____
===============
Cheers,
Stan H.